OpenStudy (anonymous):

A uniform bar is 10 ft long weighs 15 lbs. A 5 lb weight is hung at one end of the bar. At what point must the bar be supported for it to remain horizontal?

OpenStudy (anonymous):

HELP!!! @agentc0re

OpenStudy (anonymous):

HELP!!! @sami-21

OpenStudy (anonymous):

@unkabogable

OpenStudy (anonymous):

why 2.5?

OpenStudy (anonymous):

mass per feet =15/10 now 5+x1.5=1.5(10-x) where x is the distance from weight's end now 3x=10 x=3.33 no you got it??